Consider the function f(x) x3 2x4 on the interval [-2, 2] with h 0.25. Use the forward, backward, and centered finite difference approximations for the first and second derivatives so as to graphically illustrate which approximation is most accurate. Graph all three first-derivative finite difference approximations along with the theoretical, and do the same for the second derivative as well Please help me answer this question using matlab
Answer #1

=4.335 +1 8- S-8S31SCOP Calculate lheeY1ογ : = -19.131. defference la e lWe followig formula Back wand difserente - 0.2S 7.187Scalculate Ihe Cental diffevee Cent.nd d f-feve oo-드 子(2Et lrfC XED x o.2s ーチ·25 Magnitude or lte ersời =0.96577% above ssges dif-fevence al show the gaph isferene 2,l-2. L-1 f (x 2L . 3.453 125 f(Xin)= (1.75)-2+1.75 X4 5.8593 5 YWN Co. 2s)2100 16.61 /, O-l-S)121% 10. 0 Hence, the eetage from